OpenStudy (anonymous):

sec (-60) The answerr is 1 right?

jimthompson5910 (jim_thompson5910):

sec(-60) = 1/(cos(-60)) sec(-60) = 1/(cos(60)) ... because cosine is an even function sec(-60) = ???

OpenStudy (jdoe0001):

well, what does your unit circle tell you? :)

OpenStudy (anonymous):

So, following what @jim_thompson5910 said above, what is cos 60?

OpenStudy (anonymous):

cos=\[\frac{ adj }{ hyp }=\sqrt{3}\]

OpenStudy (jdoe0001):

http://i.stack.imgur.com/r8uHr.gif

OpenStudy (anonymous):

Cos 60 = 1/2 So, sec 60 = 1/cos 60 = 2

OpenStudy (anonymous):

this is same as sec(-60)

OpenStudy (anonymous):

so 60\[\sec (-60) = 1/\cos (-60) =\sqrt{3} Then 60 degree on the unit circle is \[\frac{ \pi }{ 3}\] which equals \[\left[\begin{matrix}1 & 2 \\ \sqrt{3} & 2\end{matrix}\right]\] Snices we know cos is \[\sqrt{3}\] we want the bottom number which is 2